I need ur help pls
1. The median income for full-time employed female americans is shown in the table below.
year 1980 1988 1990
$thousands 11.6 18.5 20.6
Assume t=0 corresponds to 1980, t=1 corresponds to 1981, etc.
a. Calculate the error for the best-fit linear model for this data.
b. Calculate the x-intercept for the best-fit linear model for this data.
2. What annual interest rate compounded continuously is equivalent to 4% compounded

Answer by stanbon(48568)

1. The median income for full-time employed female americans is shown in the table below.
year 1980 1988 1990
$thousands 11.6 18.5 20.6
Assume t=0 corresponds to 1980, t=1 corresponds to 1981, etc.
a. Calculate the error for the best-fit linear model for this data.
Using the TI-83 Linear Regression I get the following linear model:
y=4.5x+12.4
I don't know what you mean by the "error".
---------------
b. Calculate the x-intercept for the best-fit linear model for this data.
Let y=0, then x=-12.4/4.5=-2.75555...
------------
2. What annual interest rate compounded continuously is equivalent to 4% compounded.
Formulas:
continuous compounding for one year: A=Pe^r
4% compounding for one year: A=P(1+r)^1
e^rt=(1+0.04)^t
Take the ln of both sides to get:
rt=t[ln(1+0.04)]
rt=t[0.0392]
rate = 0.039 or 3.9%
Cheers,
Stan H.